# Don’t be lazy: CompleteP enables compute-efficient deep transformers

### v2

@@ Abstract
We study compute efficiency of LLM training when using different parameterizations, i.e., rules for adjusting model and optimizer hyperparameters (HPs) as model size changes. Some parameterizations fail to transfer optimal base HPs (such as learning rate) across changes in model depth, requiring practitioners to either re-tune these HPs as they scale up (expensive), or accept sub-optimal training when re-tuning is prohibitive. Even when they achieve HP transfer, we develop theory to show parameterizations may still exist in the lazy learning regime where layers learn only features close to their linearization, preventing effective use of depth and nonlinearity. Finally, we identify and adopt the parameterization we call CompleteP that achieves both depth-wise HP transfer and non-lazy learning in all layers. CompleteP enables a wider range of model width/depth ratios to remain compute-efficient, unlocking shapes better suited for different hardware settings and operational contexts. Moreover, CompleteP enables 12-34% compute efficiency improvements over the prior state-of-the-art.
@@ 1 Introduction
where $\mathcal{F}_{\ell}$ is the $\ell$’th residual block (for transformers, these are MLP and attention blocks). The depth-dependent re-scaling factor is governed by a single parameter $\alpha\in[0.5,1]$, where a larger $\alpha$ corresponds to stronger downscaling. Setting $\alpha=1$ gives a parameterization we call CompleteP because it is the unique $\alpha$ ensuring complete feature learning as width and depth are scaled (Section 6). Since $\alpha=0.5$ and $\alpha=1$ are the two most promising potential candidates for depth scaling [6, 7, 8], we compare these two values and find $\alpha=1$ is consistently more compute-efficient through better HP transfer and faster pre-training (Figure 1). To realize these gains, however, we must extend the parameterization to include principled re-scalings of LayerNorm (LN) and bias learning rates, AdamW’s weight decay $\lambda$, and AdamW’s $\epsilon$ (Table 1). CompleteP is simple to implement, yet yields superior upstream and downstream performance, with gains over alternative approaches increasing with model depth. To summarize, our contributions are:
@@
  • HP transfer across depth. We compare transfer of learning rate and weight initialization standard deviation across depth for the standard parameterization (SP), µP, and $\alpha=\{0.5,1\}$, when pre-training Pre-LN transformers (Figure 2, 3). We find only $\alpha=1$ enables both effective HP transfer and improved performance as model depth increases.
@@
  • Optimal shape and compute efficiency. For SP, µP, and $\alpha=\{0.5,1\}$, we construct scaling laws over models ranging from $75$M to $1.9$B total parameters, trained at a compute-optimal frontier of 20 tokens per parameter (TPP) [3]. We use this to revisit the question of compute-efficient transformer shapes (i.e. width-to-depth ratio $N\!:\!L$). The 1.9B parameter models (i.e. with 1.5B non-embedding parameters) trained with $\alpha=1$ achieve 11.8% and 34.4% FLOP savings over µP for optimally-shaped and 179-layer models, respectively (Figure 4).
@@
  • Refined desiderata for HP transfer. In Section 6, we describe the differences between SP, µP, and $\alpha\in\{0.5,1\}$. Given the strong empirical evidence for HP transfer for $\alpha=1$, we devise a refined set of desiderata for HP transfer that includes the notion of complete feature learning: we require that the learned representation in every layer of the model remains non-lazy [10] (i.e. non-linear) with respect to the parameter in both that layer and all earlier ones. Only $\alpha=1$ ensures complete feature learning, thus we name it CompleteP.
@@
  • Extra HP guidelines. In Table 1 and Appendix E, we extend existing parameterizations [8] to include prescriptions for common transformer HPs (such as weight decay, bias and LayerNorm learning rates $\eta$, and AdamW $\epsilon$) as functions of depth and width. Figure 7 shows these changes are essential for stable training with $\alpha=0.5$ in particular.
@@ 2 Related work
Early theoretical approaches for choosing HPs analyzed networks at initialization, seeking to guarantee different notions of numerical stability of the first forward and backward pass [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Subsequent work devised two parameterizations with consistent training dynamics at infinite width: the Neural Tangent Kernel (NTK) parameterization in which the model converges to its linearization around initialization [29, 30, 31] and the mean-field/µP parameterization [4, 6, 8, 10, 32, 33, 34, 35, 36, 37, 38, 39]. Training at the simultaneous limits of width and depth was explored in [40] for fully connected networks, in [6, 7] for vanilla ResNets and in [8] for transformers. Infinite-limit descriptions of training in the mean-field/µP parameterization are difficult to study analytically but give a clear proposal for HP transfer by requiring consistent dynamics of hidden layer representations across model scale [41, 42, 43, 44, 32].
This approach to HP transfer was taken up in [6, 7, 8], which derive a family of mean-field parameterizations for ResNets indexed by $\alpha\in[0.5,1]$ (see Table 1). Based on a heuristic related to feature diversity, [7] argued HP transfer was not truly possible at any $\alpha$ but that $\alpha=0.5$ was the best in practice. In contrast, we show CompleteP ($\alpha=1$) yields both HP transfer and FLOP savings during pre-training and provide theoretical justification in Section 6. Other interesting mean-field approaches to HP transfer include adaptations to sharpness-aware optimization [45], to low-precision training [46], to sparse training [47], and finally to state space models [48].
@@
Empirical approaches to HP transfer in [49, 50] normalize layer outputs and learning rates to yield HP transfer across width, consistent with the µP prescription. HP transfer across depth is often more ad hoc: [49] proposes something resembling $\alpha=0.5$, while Large et al. [50] proposes $\mathbf{h}^{\ell+1}={(\frac{L-1}{L}\cdot\mathbf{I}+L^{-1}\cdot\mathcal{F}_{\ell}(\mathbf{h}^{\ell}))}^{L}$, reminiscent of setting $\alpha=1$. Several other works, though not directly focused on HP transfer, aim to stabilize training in deep ResNets through LN modifications. Sun et al. [51] advocates multiplying pre-LN output in layer $\ell$ by $\ell^{-0.5}$, similar to taking $\alpha=0.5$. Similarly, [52] applies LN to both the input and output of each residual block, avoiding exponential-in-depth activation and gradient variance. Finally, [53] adjusts the scale of weights in each residual block, up-weights residual branch contributions, and inserts LN after the residual add. While this allows one to train very deep networks, it does not achieve depth HP transfer.
Compute-optimal transformer $N\!:\!L$ ratios.
Theoretical and empirical work have studied the optimal transformer shape. Notably Kaplan et al. [2], using SP and large-scale tests, showed a wide range of $N\!:\!L$ were close to compute-optimality, with $N\!:\!L\approx 100$ being the optimal ratio. This finding guides modern LLM shapes, which often fix $N\!:\!L\approx 100$ when parameters and tokens are scaled [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Yet SP does not fairly admit stable width and depth scaling, which undermines the prior conclusion. McLeish et al. [73] adopt a $\eta=\eta_{\text{base}}/\sqrt{L}$ parameterization resembling incomplete $\alpha=0.5$ and study compute-optimal $N\!:\!L$. However, their parameterization does not admit training stability (Figure 7) or HP transfer so their deeper models were disadvantaged, causing them to conclude shallower models are optimal. In Section 5 we revisit the transformer $N\!:\!L$ study with proper width and depth scaling control with CompleteP and show that even $N\!:\!L\approx 10$ remains close to compute optimality. Rather than fixing $N\!:\!L$, other works perform empirical searches to estimate scaling exponents for $N$ and $L$ [74, 75]. Levine et al. [76] propose a theory for transformer capacity based on separation rank, predict that optimal $N\!:\!L$ increases with parameter count, and provide empirical validation. Mixture of experts (MoE) [77] enable transformer capacity to scale without increasing depth. Similarly, [78] propose parallel sub-networks and argue that large depth isn’t necessary for competitive performance. We focus on dense transformers and consider MoE and parallel sub-networks as out of scope.
@@ 3 Methodology
We train decoder-only transformer language models [79], with pre-normalization, on the SlimPajama dataset [80]. We refer the reader to Appendix I for full methodology of all experiments. Table 1 provides an overview of how to implement the parameterizations we test in this paper. [^1] For the embedding layer, we choose an initialization variance scale independent of input dimension due to the input data being one-hot encoded. If the input data is dense, then we would require a pre-factor of $N_{\mathrm{in}}^{-1/2}$. We consider models with depth $L$ (or $2L$ residual blocks to account for both MLP and attention blocks), and width $N$ (e.g., residual activations $h^{\ell}\in\mathbb{R}^{N}$). We define adjustments in terms of width multiplier $m_{N}=N/N_{\text{base}}$ and depth multiplier $m_{L}=L/L_{\text{base}}$, where $N_{\text{base}},L_{\text{base}}$ are the width and the depth of a base model. For the base model, i.e., when $m_{N}=1,m_{L}=1$, all parameterizations are equivalent. We use $N_{\text{base}}=256,L_{\text{base}}=2$. In experiments where only $m_{N}=1$, SP is equivalent to µP; such results are labelled “SP/µP”. We extend the infinite depth parameterizations in the literature [6, 7, 8] with corrections for bias learning rate $\eta$, LayerNorm $\eta$, AdamW $\epsilon$, and weight decay $\lambda$. See Appendix E for derivations of these extensions and Appendix F for empirical verification of $\epsilon$ scaling [^2] In practice, we observe that a “small enough” $\epsilon$ is sufficient to achieve consistent dynamics and transfer across width depth. See Appendix I for details.. We provide a minimal implementation of $\alpha\in\{0.5,1.0\}$ at: https://github.com/EleutherAI/nanoGPT-mup/tree/completep.
@@ 4 Depth-wise HP transfer and $\alpha$
First, we investigate the depth-wise transfer of $\eta_{\text{base}}$ and $\sigma_{\text{base}}$ while training all models for 300M tokens with batch size $B=128$ and $\lambda_{\text{base}}=0$. We remark that this is the traditional setting where the phenomenon of HP transfer was originally observed [5]. Figure 2 shows that SP, µP, and $\alpha=0.5$ do not have stable optimal $\eta_{\text{base}}$ or $\sigma_{\text{base}}$ as depth $L$ is varied. When using the optimal HPs for $L=L_{\text{base}}=2$, the right column shows these parameterizations do not achieve consistent loss improvements with depth.
@@
For CompleteP ($\alpha=1$), the optimal $\eta_{\text{base}}$ and $\sigma_{\text{base}}$ remain stable across depths as evidenced by concentric curves with stable minima. It is the only parameterization to consistently improve loss for deeper models without HP tuning, as shown in Figure 2 right column. Such stable HP transferability dramatically reduces HP tuning budgets for deep models.
Figure 2: Depth-wise HP transfer, with 300M training tokens. Top: Learning rate ($\eta$) transfer. Bottom: Initialization standard deviation ($\sigma_{\text{init}}$) transfer. See Table 4 for experimental details.
@@
The parameterizations we test are grounded in theories derived under a fixed token count [8]. We now investigate whether the depth-wise transfer of $\eta_{\text{base}}$ extends to the compute-optimal setup prescribed in Hoffmann et al. [3], where all the models are trained for 20 TPP. We also select compute-efficient batch sizes based on total training FLOPs [2, 81, 82, 83], and ensure a well-tuned $\lambda_{\text{base}}$ [84] (see Appendix I for further details). Figure 3 shows this compute-optimal setup is much less sensitive to the choice of $\eta_{\text{base}}$ compared to training for 300M tokens with $B=128$ and $\lambda_{\text{base}}=0$ in Figure 2. Despite this reduced sensitivity, the right column shows $\alpha=1$ achieves superior losses to SP, µP, and $\alpha=0.5$ without additional $\eta_{\text{base}}$ tuning from $L_{\text{base}}$. These results lead us to conclude:
@@
Figure 3: Learning rate transfer test under compute-optimal setting (20 TPP, batch size based on BS-FLOP power law, optimal weight decay $\lambda_{\text{base}}$). See Table 4 for experimental details.

### v3

@@ Abstract
We study compute efficiency of LLM training when using different parameterizations, i.e., rules for adjusting model and optimizer hyperparameters (HPs) as model size changes. Some parameterizations fail to transfer optimal base HPs (such as learning rate) across changes in model depth, requiring practitioners to either re-tune these HPs as they scale up (expensive), or accept sub-optimal training when re-tuning is prohibitive. Even when they achieve HP transfer, we develop theory to show parameterizations may still exist in the lazy learning regime where layers learn only features close to their linearization, preventing effective use of depth and nonlinearity. Finally, we identify and adopt the parameterization we call CompleteP that achieves both depth-wise HP transfer and non-lazy learning in all layers. CompleteP enables a wider range of model width/depth ratios to remain compute-efficient, unlocking shapes better suited for different hardware settings and operational contexts. Moreover, CompleteP enables 12-34% compute efficiency improvements over the prior state-of-the-art. All experiments were run on Cerebras CS-3 systems. A minimal implementation is available at https://github.com/EleutherAI/nanoGPT-mup/tree/completep.
@@ 1 Introduction
where $\mathcal{F}_{\ell}$ is the $\ell$’th residual block (for transformers, these are MLP and attention blocks). The depth-dependent re-scaling factor is governed by a single parameter $\alpha\in[0.5,1]$.
Yang et al. [7] argue $\alpha=0.5$ works best in practice and that HP transfer is not possible at any $\alpha$, while Bordelon et al. [8] find $\alpha=1$ allows better learning as $L\to\infty$. Since $\alpha\in\{0.5,1\}$ are the two most promising potential candidates for depth scaling [6, 7, 8], we compare these two values and find $\alpha=1$ is consistently more compute-efficient through better HP transfer and faster pre-training (Figure˜1). Setting $\alpha=1$ gives a parameterization we call CompleteP because it is the unique $\alpha$ ensuring complete feature learning as width and depth are scaled (Section˜6).
@@
To realize these gains, however, we must extend the parameterization to include principled re-scalings of LayerNorm (LN) and bias learning rates, AdamW’s weight decay $\lambda$, and AdamW’s $\epsilon$ (Table 1). CompleteP is simple to implement, yet yields superior upstream and downstream performance, with gains over alternative approaches increasing with model depth. To summarize, our contributions are:
@@
  • HP transfer across depth. We compare transfer of learning rate and weight initialization standard deviation across depth (2-128 layers) for the standard parameterization (SP), µP, and $\alpha=\{0.5,1\}$, when training Pre-LN transformers (Figure˜2, 3). We find only $\alpha=1$ enables depthwise HP transfer, a feat thought impossible by Yang et al. [7].
@@
  • Optimal shape and compute efficiency. For SP, µP, and $\alpha=\{0.5,1\}$, we construct scaling laws over models ranging from 7575M to 1.91.9B total parameters, trained at a compute-optimal frontier of 20 tokens per parameter (TPP) [3]. We use this to revisit the question of compute-efficient transformer shapes (i.e. width-to-depth ratio N:LN\!:\!L). The 1.9B parameter models (i.e. with 1.5B non-embedding parameters) trained with $\alpha=1$ achieve 11.8% and 34.4% FLOP savings over µP for optimally-shaped and 179-layer models, respectively (Figure˜4).
@@
  • Refined desiderata for HP transfer. In Section˜6, we describe the differences between SP, µP, and $\alpha\in\{0.5,1\}$. Given the strong empirical evidence for HP transfer for $\alpha=1$, we devise a refined set of desiderata for HP transfer that includes the notion of complete feature learning: we require that the learned representation in every layer of the model remains non-lazy [10] (i.e. non-linear) with respect to the parameter in both that layer and all earlier ones. Only $\alpha=1$ ensures complete feature learning, thus we name it CompleteP.
@@
  • Extended parameterizations for modern training. Modern LLMs train with pre-LN and AdamW. In Table˜1 and Appendix˜D, we extend existing parameterizations [8] to include prescriptions for LayerNorm learning rates $\eta$, weight decay, bias, and AdamW $\epsilon$ as functions of depth and width. These changes are essential for stable training with $\alpha=0.5$ (Figure˜7).
@@ 2 Related work
Early methods for selecting HPs analyzed networks at initialization to ensure numerical stability in the initial forward and backward passes [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Subsequent work devised two parameterizations with consistent training dynamics at infinite width: the Neural Tangent Kernel (NTK) parameterization in which the model converges to its linearization around initialization [29, 30, 31] and the mean-field/µP parameterization [4, 6, 8, 10, 32, 33, 34, 35, 36, 37, 38, 39]. Training at the simultaneous limits of width and depth was explored in [40, 41] for fully connected networks, in [6, 7] for vanilla ResNets and in [8] for transformers. Infinite-limit descriptions of training in the mean-field/µP parameterization are difficult to study analytically but give a clear proposal for HP transfer by requiring consistent dynamics of hidden layer representations across model scale [41, 42, 43, 44, 32].
This approach to HP transfer was taken up in [6, 7, 8], which derive a family of mean-field parameterizations for ResNets indexed by $\alpha\in[0.5,1]$ (see Table˜1). Based on a heuristic related to feature diversity, [7] argued HP transfer was not truly possible at any $\alpha$ but that $\alpha=0.5$ was the best in practice. In contrast, we show CompleteP ($\alpha=1$) yields both HP transfer and FLOP savings during pre-training and provide theoretical justification in Section˜6. Other interesting mean-field approaches to HP transfer include adaptations to sharpness-aware optimization [45], to low-precision training [46], to sparse training [47], and finally to state space models [48].
@@
Empirical approaches to HP transfer in [49, 50] normalize layer outputs and learning rates to yield HP transfer across width, consistent with the µP prescription. HP transfer across depth is often more ad hoc: [49] proposes something resembling $\alpha=0.5$, while Large et al. [50] proposes $\mathbf{h}^{\ell+1}=\frac{L-1}{L}\cdot\mathbf{I}+L^{-1}\cdot\mathcal{F}_{\ell}(\mathbf{h}^{\ell})$ per layer, reminiscent of setting $\alpha=1$. Several other works, though not directly focused on HP transfer, aim to stabilize training in deep ResNets through LN modifications. Sun et al. [51] advocates multiplying pre-LN output in layer $\ell$ by $\ell^{-0.5}$, similar to taking $\alpha=0.5$. Similarly, [52] applies LN to both the input and output of each residual block, avoiding exponential-in-depth activation and gradient variance. Finally, [53] adjusts the scale of weights in each residual block, up-weights residual branch contributions, and inserts LN after the residual add. While this allows one to train very deep networks, it does not achieve depth HP transfer.
Compute-optimal transformer N:LN\!:\!L ratios.
Theoretical and empirical work have studied the optimal transformer shape. Notably Kaplan et al. [2], using SP and large-scale tests, showed a wide range of N:LN\!:\!L were close to compute-optimality, with $N\!:\!L\approx 100$ being the optimal ratio. This finding guides modern LLM shapes, which often fix $N\!:\!L\approx 100$ when parameters and tokens are scaled [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Yet SP does not fairly admit stable width and depth scaling, which undermines the prior conclusion. McLeish et al. [73] adopt a $\eta=\eta_{\text{base}}/\sqrt{L}$ parameterization resembling incomplete $\alpha=0.5$ and study compute-optimal N:LN\!:\!L. However, their parameterization does not admit training stability (Figure˜7) or HP transfer so their deeper models were disadvantaged, causing them to conclude shallower models are optimal. In Section˜5 we revisit the transformer N:LN\!:\!L study with proper width and depth scaling control with CompleteP and show that even $N\!:\!L\approx 10$ remains close to compute optimality. Rather than fixing N:LN\!:\!L, other works perform empirical searches to estimate scaling exponents for $N$ and $L$ [74, 75]. Levine et al. [76] propose a theory for transformer capacity based on separation rank, predict that optimal N:LN\!:\!L increases with parameter count, and provide empirical validation. Mixture of experts (MoE) [77] enable transformer capacity to scale without increasing depth. Similarly, [78] propose parallel sub-networks and argue that large depth isn’t necessary for competitive performance. We focus on dense transformers and consider MoE and parallel sub-networks as out of scope.
@@ 3 Methodology
For all experiments in this work, we train decoder-only Transformer language models [79] with pre-normalization, untied embeddings, ALiBi position embeddings [80] and $\text{ReLU}^{2}$ nonlinearity [81, 82], using the AdamW optimizer with decoupled weight decay [83] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e-16$. The learning rate $\eta$ schedule follows a linear warmup of min(10% of steps, 375M tokens), then a linear decay to zero [84]. We pretrain our models using an autoregressive loss (i.e. the next token prediction objective) on the SlimPajama dataset [85] with a maximal sequence length of 2048 tokens using the GPT-2 tokenizer [79]. All models use $d_{\text{head}}=64$ for each attention head and feedforward dimension $4N$. For all parameterizations, we use $\mathbf{Q}^{\top}\mathbf{K}/N$ as proposed in [5] to account for correlation between $Q$ and $K$. We use $N_{\text{base}}=256,L_{\text{base}}=2$ and scale parameters according to Table 1. All experiments were performed using Cerebras CS-3 systems. We refer the reader to Appendix˜G for full methodology of all experiments.
Table˜1 provides an overview of how to implement the parameterizations we test in this paper. [^1] For the embedding layer, we choose an initialization variance scale independent of input dimension due to the input data being one-hot encoded. If the input data is dense, then we would require a pre-factor of $N_{\mathrm{in}}^{-1/2}$. We consider models with depth $L$ (or $2L$ residual blocks to account for both MLP and attention blocks), and width $N$ (e.g., residual activations $h^{\ell}\in\mathbb{R}^{N}$). We define adjustments in terms of width multiplier $m_{N}=N/N_{\text{base}}$ and depth multiplier $m_{L}=L/L_{\text{base}}$, where $N_{\text{base}},L_{\text{base}}$ are the width and the depth of a base model. For the base model, i.e., when $m_{N}=1,m_{L}=1$, all parameterizations are equivalent. We use $N_{\text{base}}=256,L_{\text{base}}=2$. In experiments where only $m_{N}=1$, SP is equivalent to µP; such results are labelled “SP/µP”. We extend the infinite depth parameterizations in the literature [6, 7, 8] with corrections for bias learning rate $\eta$, LayerNorm $\eta$, AdamW $\epsilon$, and weight decay $\lambda$. See Appendix˜D for derivations of these extensions and practical advice for applying CompleteP to different architectures. Appendix˜E for empirical verification of $\epsilon$ scaling [^2] In practice, we observe that a “small enough” $\epsilon$ is sufficient to achieve consistent dynamics and transfer across width depth. See Appendix G for details.. We provide a minimal implementation of Table˜1 which reproduces Figure˜7 at: https://github.com/EleutherAI/nanoGPT-mup/tree/completep.
@@ 4 Depth-wise HP transfer and $\alpha$
First, we investigate the depth-wise transfer of $\eta_{\text{base}}$ and $\sigma_{\text{base}}$ while training all models for 300M tokens with batch size $B=128$ and $\lambda_{\text{base}}=0$. We remark that this is the traditional setting where the phenomenon of HP transfer was originally observed [5]. Figure˜2 shows that SP, µP, and $\alpha=0.5$ do not have stable optimal $\eta_{\text{base}}$ or $\sigma_{\text{base}}$ as depth $L$ is varied. When using the optimal HPs for $L=L_{\text{base}}=2$, the right column shows these parameterizations do not achieve consistent loss improvements with depth.
@@
For CompleteP ($\alpha=1$), the optimal $\eta_{\text{base}}$ and $\sigma_{\text{base}}$ remain stable across depths as evidenced by concentric curves with stable minima. It is the only parameterization to consistently improve loss for deeper models without HP tuning, as shown in Figure˜2 right column. Such stable HP transferability dramatically reduces HP tuning budgets for deep models. We demonstrate HP transfer from 2 to 128 layers, exceeding the depth of LLaMA-70B (80 layers) and LLaMA-405B (126 layers) [58].
Figure 2: Depth-wise HP transfer, with 300M training tokens. Top: Learning rate ($\eta$) transfer. Bottom: Initialization standard deviation ($\sigma_{\text{init}}$) transfer. See Table˜4 for experimental details.
@@
The parameterizations we test are grounded in theories derived under a fixed token count [8]. We now investigate whether the depth-wise transfer of $\eta_{\text{base}}$ extends to the compute-optimal setup prescribed in Hoffmann et al. [3], where all the models are trained for 20 TPP. We also select compute-efficient batch sizes based on total training FLOPs [2, 86, 87, 88], and ensure a well-tuned $\lambda_{\text{base}}$ [89] (see Appendix˜G for further details). Figure˜3 shows this compute-optimal setup is much less sensitive to the choice of $\eta_{\text{base}}$ compared to training for 300M tokens with $B=128$ and $\lambda_{\text{base}}=0$ in Figure˜2. Despite this reduced sensitivity, the right column shows $\alpha=1$ achieves superior losses to SP, µP, and $\alpha=0.5$ without additional $\eta_{\text{base}}$ tuning from $L_{\text{base}}$. These results lead us to conclude:
@@
Figure 3: Learning rate transfer test under compute-optimal setting (20 TPP, batch size based on BS-FLOP power law, optimal weight decay $\lambda_{\text{base}}$). See Table˜4 for experimental details.
Finding 3: Empirically, we observe larger TPP reduces HP sensitivity (Figure˜2 vs. 3).